Ashlee bought a new car. The total amount she needs to borrow is $20,839. She plans to take out a 3 year loan at an APR of 5.6%. What is her monthly payment? round to the penny.
1
Expert's answer
2020-11-10T19:07:22-0500

"interest \ per \ month = \frac{5.6} {12} = 0.467\%"


"in \ decimal form = \frac{0.467} {100} = 0.00467"


1 + 0.00467 = 1.00467


"20,839\ =\ C(1\ -\ 1\ /\ 1.00467{1.00467}^{36})\ /\ 0.00467"


"monthly \ payments\ will \ be \ \$630.2"